Myers Corporation has the following data related to direct materials costs for November: actual costs for 4,680 pounds of material, $5.00; and standard costs for 4,450 pounds of material at $6.10 per pound.

What is the direct materials price variance?

$1,403 unfavorable

$5,148 unfavorable

$5,148 favorable

$1,403 favorable

The standard costs and actual costs for direct materials for the manufacture of 2,100 actual units of product are

Standard Costs
Direct materials (per completed unit) 1,040 kilograms @$8.75

Actual Costs
Direct materials 2,100 kilograms @ $8.20

Round your final answer to the nearest dollar.

The amount of direct materials price variance is

$1,155 favorable

$572 unfavorable

$1,155 unfavorable

$572 favorable
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (6.10 - 5)*4,680

= 5178 Favourable

Option C

Materials Price Variance = (SP-AP) * AQ

= (8.75-8.20)*2,100

= 1155 Favourable

Option A